Find an equation of the tangent line to the parabola y=2x^2 at the point (-1,2), and find the x-intercept of the line. Answer by Alan3354(69443) 
 Find an equation of the tangent line to the parabola y=2x^2 at the point (-1,2), and find the x-intercept of the line.
---------------
y = 2x^2
y' = 4x
At x = -1, slope y' = -4
-------------
y = mx + b
2 = -4*(-1) + b
b = -2
-------
Ooh, x-int
y = -4x -2 is the eqn of the tangent line
0 = -4x - 2
x = -1/2 is the x-intercetp